OpenStudy (anonymous):

Please help :) Solve for x: (4x-3)/5-(2x-3)/2=-2

OpenStudy (shamil98):

\[\huge \frac{ 4x-3 }{ 5 } - \frac{ 2x-3 }{ 2 } = -2\] get a common denominator.. \[\huge \frac{ 8x-6 }{ 10 } - \frac{ 10x-15 }{ 10 } = -2\] do it from here..

OpenStudy (shamil98):

\[\huge \frac{ 8x-6-(10x-15) }{ 10 } = -2\] you have to distribute the -1 btw.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you can also multiply by 10 as a first step

OpenStudy (shamil98):

yeah, multiply 10 to everything to get rid of the fractions straight away. either way works

OpenStudy (anonymous):

yep I understand now, thank you @shamil98